OpenStudy (anonymous):

A man of mass m=70.0kg and having a density of p=1050 (kg/m^3) (while holding his breath) is completely submerged in water. (a) Write Newton's second law for this situation in term of the man's mas m, the density of water p, his volume V, and g. Neglect any viscous drag the the water. (b) Subtitle m = pV into the Newton's second law and solve for the acceleration a, canceling common factors. (c) Calculate the numeric value of the man's acceleration. (d) How long does it take the Thank You!

OpenStudy (mos1635):

 asuume positive upwords sum of forces=m*a A-W=m a p(water)*g*V-m*g=m*a p(w)*g*V-p(man)*g*V=p(man)*V*a (bye bye V) p(w)*g-p(m)*g=p(m)*a solve to a a=.......

OpenStudy (anonymous):

How do you solve for how long it will take the man to sink 8 m?

OpenStudy (mos1635):

S=1/2*a*t^2